OpenStudy (anonymous):

Ive been on this for 4 hours... F(x)= x/x-3. f(g(x))=2x. What does G(x) and how do you find it?

OpenStudy (anonymous):

now you werent given g(x) right?

OpenStudy (anonymous):

Nope. We need to find the G(x). Given the F(x) and f(g(x).

OpenStudy (anonymous):

Just to be sure... G'(x) = g(x), right? I mean, are you using G as a primitive of g?

myininaya (myininaya):

capital G? or lower case?

myininaya (myininaya):

math is case sensitive

OpenStudy (anonymous):

sorry. lowcase g

myininaya (myininaya):

\[f(g(x))=\frac{g(x)}{g(x)-3}=2x\] solve this for g: \[\frac{g}{g-3}=2x\]

OpenStudy (anonymous):

in that case its function composition

OpenStudy (anonymous):

G(x) = (-6x) / (1-2x)

myininaya (myininaya):

\[g=2x(g-3) =>g=2xg-6x=> \] \[g-2xg=-6x=>g(1-2x)=-6x=>g=\frac{-6x}{1-2x} \]

OpenStudy (anonymous):

okay, well you were given F(x) equals x/x-3 which means that F(g(x))=g(x)/g(x)-3=2x =g(x)=2x(g(x)-3) =g(x)=2x(g(x))-6x =g(x)-2x(g(x))=-6x =g(x)(1-2x)=6x =G(x)=-6x/1-2x

OpenStudy (anonymous):

\[f(g(x)) = 2x\] \[f(x) = \frac{x}{x-2} \implies f(g) = \frac{g}{g-2} \implies f(g(x)) = \frac{g(x)}{g(x) - 2} = 2x\] So, we have: \[g(x) = 2x(g(x) - 2)\] \[g(x) - 2xg(x) = -4x\] \[(1-2x)g(x) = -4x\] \[g(x) = \frac{-4x}{1-2x}\]

OpenStudy (anonymous):

Ops... I've made a mistake there. It should be x-3 on the denominator, not x-2. Sorry.

OpenStudy (anonymous):

wow thank you. Took me forever but think I get it now